OpenStudy (anonymous):

Solve for pi : C=2^pi r

OpenStudy (anonymous):

\[c=2^\pi r,2^\pi=\frac{c}{r},\pi=\frac{\ln c-\ln r}{\ln 2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2^π=c/r he just natural logged both sides then the pi comes down π* ln(2)= ln(c/r)

OpenStudy (anonymous):

So it's c over r?